# Mon CHÉRI: Mitigating Uninitialized Memory Access with Conditional Capabilities

### v2

@@ 4.3 Mon CHÉRI Support for CHERI-LLVM
LABEL:lst:ir shows an excerpt of the LLVM IR of one of the functions in LABEL:lst:attributes after CP instrumentation (see ➀ in Figure 3). The CHERI-LLVM compiler adds capability bounds to stack variables in an IR-level compiler pass, CheriBoundAllocas. This pass replaces every IR stack allocation instruction (alloca, ❶ in LABEL:lst:ir) with a llvm .cheri .cap .bounds .set intrinsic. These are then replaced with CHERI csetbounds instructions by the CHERI-RISC-V backend. We extended this pass to add llvm .cheri .cap .op .bounds .set intrinsics  (❷ inLABEL:lst:ir)for variables that are either annotated with __writebeforeread ( C in fig. 3(a)), in functions annotated with __attribute__(("writebeforeread")) ( B), or all stack variables, when compiling with the -cheri-write-before-read option ( A). The llvm .cheri .cap .op .bounds .set intrinsic is replaced with csetwbrbound instructions by the backend.
@@
The CheriBoundAllocas pass checks if alloca instructions fall within the original capability bounds and omits llvm .cheri .cap .bounds .set in those cases. We disable this optimization for variables that receive llvm .cheri .cap .op .bounds .set since it does not guarantee stores to Write-before-Read variables fall within the OBs. However, not all variables require run-time CP checks. For Write-before-Read, we optimize arrays and scalar variables allocated in function entry blocks by using a simple, non-heuristic analysis that checks if they are fully initialized before being accessed. The analysis inspects each store in the function’s first basic block and verifies whether they are preceded by loads to the same allocation. For arrays, we track initialization with a vector, checking store instructions for the base pointer or specific indices, ensuring each load is preceded by a store through dominance analysis[40]; if a load occurs before a store, the index is marked uninitialized in the vector. For scalars, we verify whether a store targets the variable, and each load is dominated by a corresponding store. Variables shown to be initialized do not receive the Write-before-Read CP and, therefore, can be checked by CheriBoundAllocas for capability-bound optimization. In § 7, we discuss the possibility of using heuristic static analysis to optimize Write-before-Read variables further.
@@
Similarly, conditional capabilities stored explicitly in memory must be updated after being used for store operations, even in straight-line code. Store linearization recursively checks if the operand used in a store is itself stored in memory. If so, an additional store operation is inserted to update the capability stored in memory. Listings LABEL:lst:64bsink and LABEL:lst:64bsinkir in § A.3 illustrate this transformation.
@@ 5 Evaluation
Following the “Good” and “Bad” classification in Juliet, Mon CHÉRI reports six false-positives. Yet, these cases do exhibit uninitialized memory access behavior (cf. § 5.1).
@@
We evaluated the Mon CHÉRI prototypes for functionality, security, performance, and area cost. The functional and security evaluation (§ 5.1) was performed on a conditional capability-enhanced QEMU-system-CHERI128 full-system emulator, CheriFreeRTOS [31], which integrates CHERI-based compartmentalization, and a Write-before-Read memory allocator (§ 4.4). For performance and area cost, we extended the CHERI-RISC-V field-programmable gate array (FPGA) softcore [44] based on the open-source Bluespec [45] RISC-V processor IP with the CP ISA extension (§ 4.1). This processor family includes Piccolo, Flute, and Tooba IP cores. Although sharing significant portions of Bluespec System Verilog (BSV) code, our performance evaluation uses the 64-bit CHERI-RISC-V Flute (RV64ACDFIMSUxCHERI). This prototype, MonCHÉRI-Flute64, was validated against RISC-V specifications using 229 RISC-V ISA tests [46].
@@ 5.1 Functional and Security Evaluation on QEMU
The tests were executed within CheriFreeRTOS, which uses an instrumented allocation API to track memory accesses. CheriFreeRTOS provides a “compartmentalize and return” mode that isolates each test case in a separate compartment, allowing test cases resulting in a CHERI protection fault to return control to the caller, which records the result and proceeds to the next test case. The test cases were compiled with the Mon CHÉRI-enhanced CHERI-LLVM at optimization level -O2 and with Write-before-Read instrumentation and store linearization but without the Write-before-Read optimization described in § 4.3.
The results in Table III show that Mon CHÉRI detects all “bad” cases (100% true-positive rate) and reports only six false positives (1% false-positive rate). Upon closer inspection, these false positives stem from cases where uninitialized variables are copied, but never used. Although these cases are technically valid violations of the Write-before-Read policy, they do not pose a security risk, as data is immediately overwritten. Analysis tools relying on taint propagation [48] might not flag these cases, as the uninitialized memory does not propagate. As Mon CHÉRI enforces policies at an architectural level, distinguishing between these benign cases and actual vulnerabilities is not currently possible. All six false positives share this pattern, where the uninitialized memory is copied but later overwritten before being used. We provide the source code for one of these cases in § A.3, LABEL:lst:juliet-cwe457-63-good. Further, all six cases have straightforward software workarounds that can be applied, once detected, to initialize variables early with a default zero value to avoid uninitialized access.
@@ 5.2 Performance and Area Evaluation on FPGA
We integrated the MonCHÉRI-Flute64 softcore into the BESSPIN-GFE security evaluation platform[49], which allows for a full-system evaluation of Mon CHÉRI performance. The GFE system includes the MonCHÉRI-Flute64 softcore, a BootROM, Soft Reset and JTAG, UART, Ethernet/DMA, DDR4, and Flash controllers. A host-based gdb debugger connects to the system over the USB/JTAG connector, and a host-based console connects over USB/UART. We measured performance using the EEMBC CoreMark [50] benchmark, running bare-metal on the GFE. Although MonCHÉRI-Flute64 supports all CPs in Table II we focus in these experiments on Write-before-Read as it is applicable to all variables in CoreMark. Consequently, applying it to all variables in the CoreMark benchmark code provides a worst-case estimate of Mon CHÉRI’s performance impact. As we expect other CPs to only be applied to a subset of variables, their impact is a fraction of Write-before-Read’s.
Table V compares the performance results MonCHÉRI-Flute64 to the CHERI-Flute64 across different configurations: no capability enforcement (no-cap), pure-capability mode (purecap), and Write-before-Read enabled in pure-capability mode (Write-before-Read + purecap ❶). The results show that the Write-before-Read extension adds a modest $\approx$ 3.5% overhead over pure-capability mode, with minimal impact on baseline performance ($\approx$ 0.4%) when capability enforcement is disabled. The combined performance impact of Write-before-Read and CHERI pure-capability mode over the baseline performance with no capability enforcement is 7%.
We also compared the performance of MonCHÉRI-Flute64 with and without store linearization (❷). Although store linearization is necessary, as explained § 5.1, for the correctness of Write-before-Read enforcement, disabling the hardware fault for this experiment allows us to compare the performance impact of the hardware changes with that of the store linearization program transformation. Enabling store linearization resulted in most of the performance degradation observed in earlier tests ($\approx$ 3.5%), with the hardware changes contributing negligible additional overhead ($\approx$ 0.2%). This suggests that performance can be improved further by optimizing the store linearization strategy.
@@ 6 Related Work
   Capstone [27] *4 ? $\approx\leavevmode\nobreak\ 50\%$5 *3
   Mon CHÉRI $\approx$ 3.5% /\ /\ / 7%6 *3
@@
    For conciseness, we include only compiler-based static analyzers focusing on uninitialized variable detection in § 6. 2 STACKLEAK protects the Linux kernel call stack after system calls.
@@
    Overhead for Mon CHÉRI given both excluding overhead for CHERI (purecap) and including overhead for CHERI, based on Table V.
Various techniques for detecting the use of uninitialized variables are routinely used in modern software development. We categorize the existing approaches into six categories, as shown in § 6. In this section, we compare conditional capabilities and Mon CHÉRI to existing approaches.
@@
Dynamic analysis, however, incurs significant performance overhead, often degrading program speed by 10$\times$ to 20$\times$, making it impractical for continuous use. These tools also struggle to accurately identify the origin of uninitialized memory. For example, Memcheck traces uninitialized variables to heap blocks or stack allocations that occur in a particular function, but may not always pinpoint the exact source. Dr. Memory, meanwhile, does not detect uninitialized variables smaller than a machine word.
@@
Sanitizers are compiler-based tools designed to detect memory-safety, concurrency, and undefined behavior issues in C and C++ programs. They intercept memory accesses via compile-time instrumentation, offering higher efficiency and accuracy than dynamic analysis tools. Sanitizers like MemorySanitizer[55] detect uninitialized stack- and heap-allocated memory at individual bit granularity, with less overhead (2$\times$ to 4$\times$ slowdown) compared to Valgrind’s one-order-of-magnitude slower dynamic analysis [76].
@@
Additionally, in Capstone, fully initialized capabilities must be explicitly promoted to regular capabilities. Either the developer or compiler must understand when a capability is expected to be fully initialized, in order to promote it. In contrast, due to the way conditional capability semantics are designed (§ 3.2), a fully initialized conditional capability is equivalent to the corresponding CHERI capability.
@@
Our evaluation demonstrates that CPs, particularly Write-before-Read capabilities in Mon CHÉRI, offer high detection accuracy with minimal false positives when used in isolation. While static analysis is valuable for early detection, tools like dynamic analysis, sanitizers, and CPs require runtime errors to be exercised. We believe that Write-before-Read CPs can complement static analysis by improving the detection of uninitialized memory issues. In § 7, we discuss how static analysis can be used with CP instrumentation to further optimize CP performance.
Our assessment of Mon CHÉRI is based on extensive evaluation using standard public sector and industry benchmarks on prototypes based on the QEMU full-system emulator and the MonCHÉRI-Flute64 softcore (§ 5). To our knowledge, no practical implementation nor compiler support for Georges et al.’s uninitialized capabilities is available to enable a fair comparison with Mon CHÉRI on FPGA. We argue that conditional capabilities, carefully designed to impose only minimal changes to the CHERI capability representation (§ 4.2), have a better chance of real-world adoption than more invasive proposals, such as Capstone.
Hardware-enforced CPs can improve detection performance for uninitialized memory issues similar to how memory tagging [9] has enhanced memory-safety sanitizers [86]. Moreover, CPs offer an alternative to automatic initialization (see § 7), emulating it without the associated drawbacks. Lastly, CPs enable novel memory access control policies, such as Write-before-Execute-Only, which provide similar benefits to memory with special-purpose features.
@@ 7 Discussion and Future Work
In § 4.3 we showed how Write-before-Read CPs can be omitted for variables that are statically verified to be initialized. Existing compiler heuristics could be used to further optimize Write-before-Read, e.g., by omitting Write-before-Read CPs when a variable is determined initialized by GCC’s -Wuninitialized, or instrument only variables identified as potentially uninitialized in certain code paths by GCC’s -Wmaybe-uninitialized. We leave further optimizations as future work as Clang/LLVM, which our conditional capability-enhanced compiler prototype is based on, does not currently replicate GCC’s -Wmaybe-uninitialized heuristic. Clang’s alternative -Wsometimes-uninitialized is more conservative as it only issues warnings when the conditions under which a variable is left uninitialized are known. Developers can already correct such cases based on the emitted warnings. We believe the Write-before-Read CPs are useful in complementing compiler-based analysis, covering cases where the analysis is inconclusive.
@@
The store linearization pass (§ 4.3) is currently limited to intra-function analysis, such as when Write-before-Read conditional capabilities are used within functions or passed to callees. However, currently propagating conditional capabilities from a callee back to the caller must be done explicitly to avoid the conditions explained in § 4.3 to occur across function boundaries. Solving inter-function store linearization is simpler in languages that enforce borrowing, where only one mutable reference to an allocation can exist at a time. Borrowing is prominent in Rust [87], but similar, compiler-enforced borrow checking has been proposed for C [88] and C++[89, 90, 91] as well. Future work should explore borrow-checker-aided full program conditional capability linearization.
@@
In § 6, we suggest Write-before-Read CPs could emulate automatic initialization, avoiding its drawbacks. Instead of issuing a hardware protection fault, a load to an address outside the conditional capability’s OB could set the destination register to zero (or a default pattern). This mimics automatic initialization without the overhead of pre-initializing memory.
Another drawback of automatic initialization, discussed in § 6, is interference with dynamic analysis and sanitizers. Initialization emulation could be controlled via a hardware configuration, allowing it to be toggled on or off. This allows the same CP-instrumented program binary to be used in production and for testing, depending on the configuration.
@@
In § 4.1 we explain how Mon CHÉRI reuses software-defined permission bits in the CHERI capability format. A drawback of the enumerator-based $p_{op}$ representation is that it makes CPs, and possible software-defined permissions, mutually exclusive.
To avoid this drawback, we considered an alternative based on that CPs always describe a subset of the conventional CHERI permissions, e.g., Write-before-Read confers the same access as R when $o$ $=$ $t$. In this alternative, a conditional control bit, $c$ is assigned from the reserved, but unused, capability bits to indicate whether the capability is in conditional mode. In conditional mode, i.e., when $c=1$, the $p_{hw}$ bits R, W, and X represent the corresponding CPs: Write-before-Read, Write-Once, and Write-before-Read respectively. This allows CP bits to overlay conventional CHERI permissions bits without losing expressivity. The Write-before-Read-Only, and Write-before-Execute-Only CPs could be overlayed similarly with an additional exclusive permission bit.
@@
In § 4.1, we assume that memory accessible via CPs is written sequentially. This holds for most data types and operations like memcpy() and memset(), but not for C structures that are initialized field-by-field or contain uninitialized padding. Software workarounds like #pragma pack with ordered field-initialization or an initial memset() before individual fields are set can prevent false positives from field-by-field access.
@@
In § 4.2, we explain that the current operation bounds encoding restricts conditional capabilities to $I_{E}\leq 2$. Similar to how the CHERI Concentrate encoding sacrifices alignment precision as allocation sizes increase, the operation top ($o$) can sacrifice write precision as $I_{E}$ increases. At an architectural level, this is achieved by zero-padding the least significant bits of $O_{E}$, akin to handling $B$ and $T$. When the precision of the least significant stored bit in $O_{E}$ exceeds the ability to express writes to individual bytes, halfwords, or words, corresponding store instructions (s[bhw]) are disabled for that conditional capability. Extending writes beyond doubleword precision requires a variant of SetOpBounds, which extends $o$ under the condition that the two preceding instructions have been writes reaching a target granularity.

### v3

@@ 4.3 Mon CHÉRI Support for CHERI-LLVM
LABEL:lst:ir shows an excerpt of the LLVM IR of one of the functions in LABEL:lst:attributes after CP instrumentation (see ➀ in Figure˜3). The CHERI-LLVM compiler adds capability bounds to stack variables in an IR-level compiler pass, CheriBoundAllocas. This pass replaces every IR stack allocation instruction (alloca, ❶ in LABEL:lst:ir) with a llvm .cheri .cap .bounds .set intrinsic. These are then replaced with CHERI csetbounds instructions by the CHERI-RISC-V backend. We extended this pass to add llvm .cheri .cap .op .bounds .set intrinsics (❷ inLABEL:lst:ir)for variables that are either annotated with __writebeforeread ( C in fig.˜3(a)), in functions annotated with __attribute__(("writebeforeread")) ( B), or all stack variables, when compiling with the -cheri-write-before-read option ( A). The llvm .cheri .cap .op .bounds .set intrinsic is replaced with csetwbrbound instructions by the backend.
@@
The CheriBoundAllocas pass checks if alloca instructions fall within the original capability bounds and omits llvm .cheri .cap .bounds .set in those cases. We disable this optimization for variables that receive llvm .cheri .cap .op .bounds .set since it does not guarantee stores to Write-before-Read variables fall within the OBs. However, not all variables require run-time CP checks. For Write-before-Read, we optimize arrays and scalar variables allocated in function entry blocks by using a simple, non-heuristic analysis that checks if they are fully initialized before being accessed. The analysis inspects each store in the function’s first basic block and verifies whether they are preceded by loads to the same allocation. For arrays, we track initialization with a vector, checking store instructions for the base pointer or specific indices, ensuring each load is preceded by a store through dominance analysis[40]; if a load occurs before a store, the index is marked uninitialized in the vector. For scalars, we verify whether a store targets the variable, and each load is dominated by a corresponding store. Variables shown to be initialized do not receive the Write-before-Read CP and, therefore, can be checked by CheriBoundAllocas for capability-bound optimization. In §˜7, we discuss the possibility of using heuristic static analysis to optimize Write-before-Read variables further.
@@
Similarly, conditional capabilities stored explicitly in memory must be updated after being used for store operations, even in straight-line code. Store linearization recursively checks if the operand used in a store is itself stored in memory. If so, an additional store operation is inserted to update the capability stored in memory. Listings LABEL:lst:64bsink and LABEL:lst:64bsinkir in §˜A.3 illustrate this transformation.
@@ 5 Evaluation
Following the “Good” and “Bad” classification in Juliet, Mon CHÉRI reports six false-positives. Yet, these cases do exhibit uninitialized memory access behavior (cf. §˜5.1).
@@
We evaluated the Mon CHÉRI prototypes for functionality, security, performance, and area cost. The functional and security evaluation (§˜5.1) was performed on a conditional capability-enhanced QEMU-system-CHERI128 full-system emulator, CheriFreeRTOS [31], which integrates CHERI-based compartmentalization, and a Write-before-Read memory allocator (§˜4.4). For performance and area cost, we extended the CHERI-RISC-V field-programmable gate array (FPGA) softcore [44] based on the open-source Bluespec [45] RISC-V processor IP with the CP ISA extension (§˜4.1). This processor family includes Piccolo, Flute, and Tooba IP cores. Although sharing significant portions of Bluespec System Verilog (BSV) code, our performance evaluation uses the 64-bit CHERI-RISC-V Flute (RV64ACDFIMSUxCHERI). This prototype, MonCHÉRI-Flute64, was validated against RISC-V specifications using 229 RISC-V ISA tests [46].
@@ 5.1 Functional and Security Evaluation on QEMU
The tests were executed within CheriFreeRTOS, which uses an instrumented allocation API to track memory accesses. CheriFreeRTOS provides a “compartmentalize and return” mode that isolates each test case in a separate compartment, allowing test cases resulting in a CHERI protection fault to return control to the caller, which records the result and proceeds to the next test case. The test cases were compiled with the Mon CHÉRI-enhanced CHERI-LLVM at optimization level -O2 and with Write-before-Read instrumentation and store linearization but without the Write-before-Read optimization described in §˜4.3.
The results in Table˜III show that Mon CHÉRI detects all “bad” cases (100% true-positive rate) and reports only six false positives (1% false-positive rate). Upon closer inspection, these false positives stem from cases where uninitialized variables are copied, but never used. Although these cases are technically valid violations of the Write-before-Read policy, they do not pose a security risk, as data is immediately overwritten. Analysis tools relying on taint propagation [48] might not flag these cases, as the uninitialized memory does not propagate. As Mon CHÉRI enforces policies at an architectural level, distinguishing between these benign cases and actual vulnerabilities is not currently possible. All six false positives share this pattern, where the uninitialized memory is copied but later overwritten before being used. We provide the source code for one of these cases in §˜A.3, LABEL:lst:juliet-cwe457-63-good. Further, all six cases have straightforward software workarounds that can be applied, once detected, to initialize variables early with a default zero value to avoid uninitialized access.
@@ 5.2 Performance and Area Evaluation on FPGA
We integrated the MonCHÉRI-Flute64 softcore into the BESSPIN-GFE security evaluation platform[49], which allows for a full-system evaluation of Mon CHÉRI performance. The GFE system includes the MonCHÉRI-Flute64 softcore, a BootROM, Soft Reset and JTAG, UART, Ethernet/DMA, DDR4, and Flash controllers. A host-based gdb debugger connects to the system over the USB/JTAG connector, and a host-based console connects over USB/UART. We measured performance using the EEMBC CoreMark [50] benchmark, running bare-metal on the GFE. Although MonCHÉRI-Flute64 supports all CPs in Table˜II we focus in these experiments on Write-before-Read as it is applicable to all variables in CoreMark. Consequently, applying it to all variables in the CoreMark benchmark code provides a worst-case estimate of Mon CHÉRI’s performance impact. As we expect other CPs to only be applied to a subset of variables, their impact is a fraction of Write-before-Read’s.
Table˜V compares the performance results MonCHÉRI-Flute64 to the CHERI-Flute64 across different configurations: no capability enforcement (no-cap), pure-capability mode (purecap), and Write-before-Read enabled in pure-capability mode (Write-before-Read + purecap ❶). The results show that the Write-before-Read extension adds a modest $\approx$ 3.5% overhead over pure-capability mode, with minimal impact on baseline performance ($\approx$ 0.4%) when capability enforcement is disabled. The combined performance impact of Write-before-Read and CHERI pure-capability mode over the baseline performance with no capability enforcement is 7%.
We also compared the performance of MonCHÉRI-Flute64 with and without store linearization (❷). Although store linearization is necessary, as explained §˜5.1, for the correctness of Write-before-Read enforcement, disabling the hardware fault for this experiment allows us to compare the performance impact of the hardware changes with that of the store linearization program transformation. Enabling store linearization resulted in most of the performance degradation observed in earlier tests ($\approx$ 3.5%), with the hardware changes contributing negligible additional overhead ($\approx$ 0.2%). This suggests that performance can be improved further by optimizing the store linearization strategy.
@@ 6 Related Work
   Capstone [27] *4 ? $\approx\penalty 10000\ 50\%$5 *3
   Mon CHÉRI $\approx$ 3.5% /\ /\ 7%6 *3
@@
    For conciseness, we include only compiler-based static analyzers focusing on uninitialized variable detection in Table˜VI. 2 STACKLEAK protects the Linux kernel call stack after system calls.
@@
    Overhead for Mon CHÉRI given both excluding overhead for CHERI (purecap) and including overhead for CHERI, based on Table˜V.
Various techniques for detecting the use of uninitialized variables are routinely used in modern software development. We categorize the existing approaches into six categories, as shown in Table˜VI. In this section, we compare conditional capabilities and Mon CHÉRI to existing approaches.
@@
[…]
@@
[…]
@@
Additionally, in Capstone, fully initialized capabilities must be explicitly promoted to regular capabilities. Either the developer or compiler must understand when a capability is expected to be fully initialized, in order to promote it. In contrast, due to the way conditional capability semantics are designed (§˜3.2), a fully initialized conditional capability is equivalent to the corresponding CHERI capability.
@@
Our evaluation demonstrates that CPs, particularly Write-before-Read capabilities in Mon CHÉRI, offer high detection accuracy with minimal false positives when used in isolation. While static analysis is valuable for early detection, tools like dynamic analysis, sanitizers, and CPs require runtime errors to be exercised. We believe that Write-before-Read CPs can complement static analysis by improving the detection of uninitialized memory issues. In §˜7, we discuss how static analysis can be used with CP instrumentation to further optimize CP performance.
Our assessment of Mon CHÉRI is based on extensive evaluation using standard public sector and industry benchmarks on prototypes based on the QEMU full-system emulator and the MonCHÉRI-Flute64 softcore (§˜5). To our knowledge, no practical implementation nor compiler support for Georges et al.’s uninitialized capabilities is available to enable a fair comparison with Mon CHÉRI on FPGA. We argue that conditional capabilities, carefully designed to impose only minimal changes to the CHERI capability representation (§˜4.2), have a better chance of real-world adoption than more invasive proposals, such as Capstone.
Hardware-enforced CPs can improve detection performance for uninitialized memory issues similar to how memory tagging [9] has enhanced memory-safety sanitizers [86]. Moreover, CPs offer an alternative to automatic initialization (see §˜7), emulating it without the associated drawbacks. Lastly, CPs enable novel memory access control policies, such as Write-before-Execute-Only, which provide similar benefits to memory with special-purpose features.
@@ 7 Discussion and Future Work
In §˜4.3 we showed how Write-before-Read CPs can be omitted for variables that are statically verified to be initialized. Existing compiler heuristics could be used to further optimize Write-before-Read, e.g., by omitting Write-before-Read CPs when a variable is determined initialized by GCC’s -Wuninitialized, or instrument only variables identified as potentially uninitialized in certain code paths by GCC’s -Wmaybe-uninitialized. We leave further optimizations as future work as Clang/LLVM, which our conditional capability-enhanced compiler prototype is based on, does not currently replicate GCC’s -Wmaybe-uninitialized heuristic. Clang’s alternative -Wsometimes-uninitialized is more conservative as it only issues warnings when the conditions under which a variable is left uninitialized are known. Developers can already correct such cases based on the emitted warnings. We believe the Write-before-Read CPs are useful in complementing compiler-based analysis, covering cases where the analysis is inconclusive.
@@
The store linearization pass (§˜4.3) is currently limited to intra-function analysis, such as when Write-before-Read conditional capabilities are used within functions or passed to callees. However, currently propagating conditional capabilities from a callee back to the caller must be done explicitly to avoid the conditions explained in §˜4.3 to occur across function boundaries. Solving inter-function store linearization is simpler in languages that enforce borrowing, where only one mutable reference to an allocation can exist at a time. Borrowing is prominent in Rust [87], but similar, compiler-enforced borrow checking has been proposed for C [88] and C++[89, 90, 91] as well. Future work should explore borrow-checker-aided full program conditional capability linearization.
@@
In §˜6, we suggest Write-before-Read CPs could emulate automatic initialization, avoiding its drawbacks. Instead of issuing a hardware protection fault, a load to an address outside the conditional capability’s OB could set the destination register to zero (or a default pattern). This mimics automatic initialization without the overhead of pre-initializing memory.
Another drawback of automatic initialization, discussed in §˜6, is interference with dynamic analysis and sanitizers. Initialization emulation could be controlled via a hardware configuration, allowing it to be toggled on or off. This allows the same CP-instrumented program binary to be used in production and for testing, depending on the configuration.
@@
In §˜4.1 we explain how Mon CHÉRI reuses software-defined permission bits in the CHERI capability format. A drawback of the enumerator-based $p_{op}$ representation is that it makes CPs, and possible software-defined permissions, mutually exclusive.
To avoid this drawback, we considered an alternative based on that CPs always describe a subset of the conventional CHERI permissions, e.g., Write-before-Read confers the same access as R when $o$ == $t$. In this alternative, a conditional control bit, $c$ is assigned from the reserved, but unused, capability bits to indicate whether the capability is in conditional mode. In conditional mode, i.e., when $c=1$, the $p_{hw}$ bits R, W, and X represent the corresponding CPs: Write-before-Read, Write-Once, and Write-before-Read respectively. This allows CP bits to overlay conventional CHERI permissions bits without losing expressivity. The Write-before-Read-Only, and Write-before-Execute-Only CPs could be overlayed similarly with an additional exclusive permission bit.
@@
In §˜4.1, we assume that memory accessible via CPs is written sequentially. This holds for most data types and operations like memcpy() and memset(), but not for C structures that are initialized field-by-field or contain uninitialized padding. Software workarounds like #pragma pack with ordered field-initialization or an initial memset() before individual fields are set can prevent false positives from field-by-field access.
@@
In §˜4.2, we explain that the current operation bounds encoding restricts conditional capabilities to $I_{E}\leq 2$. Similar to how the CHERI Concentrate encoding sacrifices alignment precision as allocation sizes increase, the operation top ($o$) can sacrifice write precision as $I_{E}$ increases. At an architectural level, this is achieved by zero-padding the least significant bits of $O_{E}$, akin to handling $B$ and $T$. When the precision of the least significant stored bit in $O_{E}$ exceeds the ability to express writes to individual bytes, halfwords, or words, corresponding store instructions (s[bhw]) are disabled for that conditional capability. Extending writes beyond doubleword precision requires a variant of SetOpBounds, which extends $o$ under the condition that the two preceding instructions have been writes reaching a target granularity.